17 students after 6 weeks mean is 84, SD= 16 N=17. Previous mean=78. Is difference consistent
1
Expert's answer
2019-07-24T13:44:25-0400

One sample t-test.

Null hypothesis "H_0:\;\mu=78"

Alternative hypothesis "H_a:\;\mu\neq78"

Test statistic: "t=\frac{\bar{x}-\mu}{\frac{s}{\sqrt{n}}}=\frac{84-78}{\frac{16}{\sqrt{17}}}=1.55"

Degrees of freedom: "df=n-1=17-1=16"

P-value: "p=0.1407"

Since the P-value is greater than 0.05, fail to reject the null hypothesis.

The difference is not consistent..